### v2

# Lodz, Pomorska 149/153, 90-236, Poland

Abstract
We study the area and volume laws for entanglement of free quantum scalar fields. In addition to the entropy, we use the notion of the capacity of entanglement, which measures entropy fluctuations. We consider flat spacetimes as well as the curved ones relevant for cosmology. Moreover, we put special emphasis on quench phenomena and different geometries of the entangling surfaces.
First, we show that, in the Minkowski spacetime, the capacity of entanglement, like entropy, exhibits the area law for two kinds of geometries of the entangling surfaces: the sphere and strip. Moreover, we show that the ratio of both quantities takes the same values for both surfaces. Next, we turn our attention to quenches. Namely, we analyse the dynamics of capacity; in particular, contribution of the volume and surface terms. Moreover, we compare these results with theoretical predictions resulting from the quasiparticles model. In the second part, we consider the above issues for the FLRW spaces; especially, for de Sitter space as well as a metric modeling the transition to radiation-dominated era. Finally, we analyse the abrupt quenches in de Sitter space.
1 Introduction
Quantum entanglement is one of the fascinating aspects of quantum systems, and its implications go beyond quantum information processing and technology. For example, it seems to play the relevant role in the statistical mechanics, condensed matter or high energy physics, see b1a -b1e for wider discussions and further references. The latter aspect is particularly interesting because of quantum gravitational effects. As a result, the notion of entanglement in quantum field theory has been extensively investigated in the recent years (especially for the low-dimensional or flat spacetimes). An important example of such studies is the concept of entropy in black hole physics; in particular, its relation to the area instead of volume scaling, see the pioneering works b4a ; b4b ; b4c and b3a ; b3b for review, or more generally the notion of the holographic entanglement entropy b5a ; b5b .
To quantify entanglement several measures have been proposed. The most popular one, for a pure state of a bipartite system, is the von Neumann entropy. One of the interesting properties of the von Neumann entropy is that it can be written as the expectation value of the so-called modular Hamiltonian (i.e. the minus logarithm of the density operator). In view of this, the variance of the modular Hamiltonian can serve as a measure of fluctuations of the entanglement entropy b7 . On the other hand, it turns out that this variance can be treated as a kind of “heat" capacity. Such an approach originally appeared in the context of condensed mater physics b8 . However, recently such a (modular) capacity gained some additional attention due to quantum gravitational effects, holographic duality, and other aspects of the field theory, see e.g. b6a -b6o .
Motivated by these results, in this work we continue the study of the entropy and capacity of entanglement, however, with the emphasis on higher dimensional (in particular (1+3)-dimensional) flat and non-flat spacetimes as well as various geometries of the entangling surfaces. Moreover, we put special attention on the dynamics of these quantities during quenches; in particular, the analysis of the area and volume laws. Such investigations are relevant for various physical processes, including thermalization or non-equilibrium systems, see e.g. b12a -b12g .
The work is organized as follows. In Sec. 2 we recall the discretization procedure for fields in non-stationary spacetimes as well as the formalism needded to compute the entropy and capacity in this approach. In Sec. 3 we investigate the capacity for $(1+3)$ and $(1+2)$-dimensional Minkowski spacetimes and the spherical entangling surface; we consider the case of constant mass as well as mass quenches. In the latter case we compare the numerical results with the quasiparticles model (basing on EPR pairs). To analyse universal features of the quasiparticles model as well as of the ratio of the capacity and entropy in Sec. 4 we make analogous considerateness for the strip geometry. In Sec. 5 we consider curved spacetimes. We focus on the FLRW metrics because of their cosmological applications and holographic aspects. In particular, we investigate the above issues for de Sitter (dS) space as well as after transition to the radiation-dominated era. Finally, we consider quenches in dS space. The conclusions are collected in Sec. 6, while some technical details and auxiliary facts are provided in Appendices A and B.
2 Setup
Let us consider the free scalar field $\Phi$ in the curved spacetime $g_{\mu\nu}$, described by the action
$\frac{1}{2}\int d^{4}x\sqrt{-g}(g^{\mu\nu}\partial_{\mu}\Phi\partial_{\nu}\Phi+m^{2}(x^{0})\Phi^{2}),$ (1)
where $m(x^{0})$ is, in general, a “time-dependent" mass parameter modeling the quench. Obviously, the simplest and most relevant case is the Minkowski spacetime. However, other spacetimes can be also considered; for example, the FLRW metrics are of great interest, due to their applications in cosmology, or more generally time-dependent spherically symmetric spacetimes. For such spaces we can perform the discretization procedure of the action (1), see e.g. b13a ; b13b ; b13c . As a result we get the Hamiltonian on the lattice [^1] Throughout our considerations we put the lattice spacing equal one. (together with appropriate boundary conditions). Namely, in $1+1$ dimensions the resulting Hamiltonian can be written in the form
$H(x_{0})=\frac{1}{2}\sum_{j=1}^{N}\pi_{j}^{2}+\frac{1}{2}\phi^{T}\Lambda(x_{0})\phi,$ (2)
where $\phi=(\phi_{1},\ldots,\phi_{N})$ and $\Lambda(x_{0})$ is a symmetric $N\times N$ matrix build with suitable frequencies and couplings parameters; in higher dimensions the Hamiltonian is the sum of the ones given in eq. (2), see e.g. b13a . In view of the above, the discretization procedure enables to analyse various aspects of the field theory and even for constant mass leads to interesting issues. One of them is the meaning of entanglement in quantum field theory. In this case, we divide the space in two regions. In the lattice approach, this corresponds to a splinting of the whole system into two parts (consisting with $n$ and NnN-n oscillators, respectively). Next, we define the reduced density operator with respect to the one part.
To quantify quantum correlations various measures have been proposed. The von Neumann entropy, or more generally, the Rényi entropy $R_{\alpha}$ are the most common. It is worth noting that the von Neumann entropy can be written as the expectation value of the operator $K=-\ln(\rho)$, the so-called modular Hamiltonian. In view of this the variance of $K$, i.e. $C=\langle K^{2}\rangle-\langle K\rangle^{2}$, can be considered as a measure of the fluctuations of the entanglement entropy. On the hand, following a thermodynamical analogy, see e.g. Refs. b8 ; b6a ; boer , $C$ can be treated as the “heat" capacity; this in turn leads to an equivalent definition and terminology for $C$, namely the (modular) capacity:
$C=(\partial_{\alpha}^{2}((1-\alpha)R_{\alpha}))|_{\alpha=1}\ .$ (3)
In this work we will study the entanglement entropy and its fluctuation, i.e. capacity of entanglement, for (quenched) fields in some (curved) spacetimes. In view of the previous discussion concerning the lattice procedure these problems can be reduced to the ones for the discretized systems. Thus, first, we briefly recall the main steps of such an approach b6l ; b10a ; b15 . Namely, we start with the instantaneous ground state (at some initial time, subscript 0 in the notation) of the whole system. Then, the evolution of the density matrix [^2] For simplicity of notation we omit the time parameter $t$ in the matrices and subsequent quantities and dot refers to derivative with respect to $t$. is given by
$\rho(\phi,\phi^{\prime})=\sqrt{\det(\Omega/\pi)}\exp(i\phi^{T}B\phi-i{\phi^{\prime}}^{T}B\phi^{\prime}-\frac{1}{2}\phi^{T}\Omega\phi-\frac{1}{2}{\phi^{\prime}}^{T}\Omega\phi^{\prime}),$ (4)
$\Omega=U^{T}\sqrt{\tilde{\Lambda}}U$, $B=U^{T}\tilde{B}U$ where $\tilde{B},\tilde{\Lambda}$ are diagonal matrices with elements $(\tilde{\Lambda})_{ij}=\lambda_{i}^{0}/b_{i}^{4}\delta_{ij}$ and $(\tilde{B})_{ij}=\dot{b}_{i}/(2b_{i})\delta_{ij}$, respectively, while $b_{j}$ are the solutions of the Ermakov equations with the frequencies $\lambda_{j}$
$\overset{..}{b_{j}}+\lambda_{j}b_{j}=\frac{\lambda_{j}^{0}}{b^{3}_{j}},\quad j=1,\ldots,N;$ (5)
and, finally, $U$ is a time-independent matrix diagonalizing $\Lambda$, i.e. $U\Lambda U^{T}=Diag(\lambda_{1},\ldots,\lambda_{N}).$ Next, we split the whole system into two parts: the first one $\mathcal{A}$ consisting of the first $n$ oscillators and the second one $\mathcal{B}$ related to the remaining NnN-n ones. To find the reduced density we rewrite $\Omega$ and $B$ in the form
$\Omega=\left(\begin{array}[]{cc}\Omega_{1}&\Omega_{2}\\ \Omega_{2}^{T}&\Omega_{3}\end{array}\right),\quad B=\left(\begin{array}[]{cc}B_{1}&B_{2}\\ B_{2}^{T}&B_{3}\end{array}\right),$ (6)
where $\Omega_{1},B_{1}$ are $n\times n$ matrices. Next, integrating over the subsystem $\mathcal{A}$ we get the reduced density of the subsystem $\mathcal{B}$}
$\rho_{\mathcal{B}}(\phi_{\mathcal{B}},\phi_{\mathcal{B}}^{\prime})=A\exp(i\phi_{\mathcal{B}}^{T}Z\phi_{\mathcal{B}}-i{\phi_{\mathcal{B}}^{\prime}}^{T}Z\phi_{\mathcal{B}}^{\prime}-\frac{1}{2}\phi_{\mathcal{B}}^{T}\Upsilon\phi_{\mathcal{B}}-\frac{1}{2}{\phi_{\mathcal{B}}^{\prime}}^{T}\Upsilon\phi_{\mathcal{B}}^{\prime}+\phi_{\mathcal{B}}^{T}\Delta\phi_{\mathcal{B}}^{\prime}),$ (7)
where $\phi_{\mathcal{B}}=(\phi_{n+1}\ldots,\phi_{N})$ and $Z,\Upsilon,\Delta$ are $(N-n)\times(N-n)$ matrices given by
$Z$ $=B_{3}-B_{2}^{T}\Omega_{1}^{-1}\Omega_{2},$ (8)
$\Upsilon$ $=\Omega_{3}-\frac{1}{2}\Omega_{2}^{T}\Omega_{1}^{-1}\Omega_{2}+2B_{2}^{T}\Omega_{1}^{-1}B_{2},$ (9)
$\Delta$ $=\frac{1}{2}\Omega_{2}^{T}\Omega_{1}^{-1}\Omega_{2}+2B_{2}^{T}\Omega_{1}^{-1}B_{2}+i\Theta,$ (10)
with $\Theta=\Omega_{2}^{T}\Omega_{1}^{-1}B_{2}-B_{2}^{T}\Omega_{1}^{-1}\Omega_{2}$. The spectrum of the density operator with the Hermitian matrix $\Delta$ was discussed, in Ref. b15 . It turns out that it is of the form
$(1-\xi_{1})(1-\xi_{2})\ldots(1-\xi_{N-n})\xi_{1}^{m_{1}}\xi_{2}^{m_{2}}\ldots\xi_{N-n}^{m_{N-n}},$ (11)
where $\xi$’s are the inverse of the eigenvalues (larger than one) of the following matrix
$\left(\begin{array}[]{cc}2\tilde{\Delta}^{-1}&-\tilde{\Delta}^{-1}\tilde{\Delta}^{T}\\ I&0\end{array}\right),$ (12)
where
$\tilde{\Delta}=(\tilde{\Upsilon})^{-1/2}\Pi\Delta\Pi^{T}(\tilde{\Upsilon})^{-1/2},$ (13)
while $\Pi$ is an orthogonal matrix diagonalizing $\Upsilon$, i.e. $\Pi\Upsilon\Pi^{T}=\tilde{\Upsilon}$. Now, from (11), definition of the von Neumann entropy and eq. (3) we get
$S=-\sum_{j=n+1}^{N}\left(\ln(1-\xi_{j})+\frac{\xi_{j}}{1-\xi_{j}}\ln(\xi_{j})\right),\quad C=\sum_{j=n+1}^{N}\frac{\xi_{j}\ln^{2}(\xi_{j})}{(1-\xi_{j})^{2}}.$ (14)
As noted above, in higher dimensional spacetimes the total Hamiltonian is a sum of independent Hamiltonians of the form (2) (see, e.g., b13a ), thus entropy and capacity are also the sum of the corresponding components. For example, in $1+3$ dimensions the discretization procedure yields $S=\sum_{lm}S^{lm}$ and $C=\sum_{lm}C^{lm}$ where $S^{lm}$ and $C^{lm}$ are of the form as above, see Appendix A for more details.
Alternatively, to obtain the entropy and capacity, we can use another approach based on the correlation (covariance) matrix and symplectic spectrum. This approach seems more useful for numerical computations and has been successfully applied in various investigations, more details can be found, e.g., in Refs. cotler ; b12g . Therefore, we also apply this approach to numerical calculations. However, to obtain some analytical results we will employ the aforementioned approach based on the eigenvalues of the reduced density. Finally, let us note that the relation between both the approaches has been recently discussed in more detail in Ref. b16 .
3 Minkowski spacetime - spherical geometry
3.1 $(1+3)$ dimensions
A remarkable property of the (geometric) entanglement entropy of the scalar field with constant mass (at its ground state) is that it obeys the area law, i.e. at the leading order the entropy is proportional to the area of the entangling sphere $S=a_{1}R^{2}$ b4a ; b4b ; b4c ; in the notation of Sec. 2, $R=n+1/2$. In turns out that the same situation holds for the capacity. In fact, plotting the capacity [^3] In our considerations we take $l$’s so large that the entropy and capacity do not change significantly, i.e. a few thousand; this is consistent with other considerations reported in the literature, see e.g. b6n and references therein. as the function of $R^{2}$ for the ground state of the discretized Hamiltonian, see Sec. 2 and Appendix A, we observe, in the left part of Fig. 1, that $C=a_{2}R^{2}$. However, let us stress that the slopes $a_{1},a_{2}$ for the entropy and capacity, respectively, are different and both depend on mass. In view of this the ratio $C/S$, at the leading order, does not depend on $R$ ($C/S=a_{2}/a_{1}$). Now, let us consider this ratio for various values of the mass parameter, see the right panel in Fig. 1. For the massless field the slope of capacity is $a_{2}\simeq 1.56$ while for entropy $a_{1}\simeq 0.3$, thus the ratio $C/S$ is approximately equal to $5.2$ what coincides with the results of Ref. boer ; next, it increases with mass, see Fig. 1. In general, this ratio is scheme dependent; though, for some conformal theories (with a dual holographic gravity description) is universal and equal to one, see Ref. boer for a more extensive discussion on this subject. Here, we will analyse another aspect of $C/S$. Namely, the dependence on the geometry of the entangling surface. To this end in Sec. 4 we will make a similar analysis for the strip geometry and show that both cases coincide very well; this, in turn, suggests another universal property of this ratio.
Figure 1: The $(1+3)$-dimensional Minkowski spacetime - the spherical geometry. The left panel: $m=0$, entropy (blue data points) and capacity (orange data points) with respect to $R^{2}$. The right panel: the ratio $C/S=a_{2}/a_{1}$ with respect to $m$, for $m=0$ it is $5.2$.
Now, let us consider a more complicated situation when mass of the field is time-dependent. We will analyse the abrupt quench; though the continuous protocols (e.g. related to hyperbolic tangent) can be also considered. Moreover, we will focus on the case where the final mass is equal to zero (the critical protocol). So, we start with a field of mass $m_{i}$ and next at time $t_{0}=0$ there is a sudden change of mass to zero. In this case, the solutions of the Ermakov equations are given by the formulae
$b_{j}(t)=\sqrt{r_{j}\cos(2t\sqrt{\lambda_{j}(f)})+s_{j}},$ (15)
where $r_{j}=(\lambda_{j}(f)-\lambda_{j}(i))/(2\lambda_{j}(f))$, $s_{j}=(\lambda_{j}(f)+\lambda_{j}(i))/(2\lambda_{j}(f))$, and $\lambda_{j}(i),\lambda_{j}(f)$ are the eigenvalues of $\Lambda$ before and after the quench, respectively. We begin our analysis with the temporal evolution of the capacity for various $R$ (equivalently $n$). The typical dynamics of the entropy and capacity is presented in Fig. 2. Let us note that their values grow with $n$; namely, for $n<N/2$ they are smaller than for $n=N/2$. Moreover, we see that the dynamics of the entropy and capacity have an increasing period, up to $t=R=(n+1/2)$, and finally they oscillate around some asymptotic values (depending on $n$).
Figure 2: The abrupt quench in the $(1+3)$-dimensional Minkowski space - spherical geometry; $m_{i}=10$, $m_{f}=0$, $N=60$. The left panel: entropy - blue $n=15$, green $n=25$, black $n=30$. The right panel: capacity - yellow $n=15$, orange $n=25$, red $n=30$.
To analyse the area law, let us make, for fixed time $t$, the decomposition
$S(t)=a_{1}(t)R^{2}+b_{1}(t)R^{3},\quad C(t)=a_{2}(t)R^{2}+b_{2}(t)R^{3}.$ (16)
Then the coefficients $a_{i}(t)$ and $b_{i}(t)$, for $i=1,2$, describe the impact of the surface and volume terms, respectively. In particular, when $Rb_{i}(t)/a_{i}(t)\ll 1$ the are law holds, in the other case the volume term has to be taken into account. To analyse the dynamics of the area law we plot the ratio $b_{i}/a_{i}$ ($i=1$ - entropy, $i=2$ - capacity), see Fig. 3.
Figure 3: The ratio $b_{i}/a_{i}$ for the abrupt quench in the $(1+3)$-dimensional Minkowski space - spherical geometry; $N=60$, $m_{f}=0$. Blue curve - $b_{1}/a_{1}$ for entropy, red curve - $b_{2}/a_{2}$ for capacity. The left panel: $m_{i}=10$. The right panel: $m_{i}=0.5$.
Obviously, the area law holds before the quench (in our case for $t\leq 0$). However, for sufficient small times the surface term is also dominant; in fact, for $t=2$ and even for relatively large gap, i.e. $m_{i}=10$, the values of the entropy and capacity with respect to $R^{2}$ fit quite well to the straight lines, see the left panel of Fig. 4. Now, let us focus on large times. Then the situation depends on the initial mass. When the gap is large (e.g. $m_{i}=10$) then the volume term remains decisive; for smaller gap the mixture of the volume and surface terms is necessary. Indeed, for a large initial mass the ratio $b_{i}(t)/a_{i}(t)$ oscillate about some asymptotic values, see also Fig. 7 in Sec. 3.2. Finally, let us not that the volume term is particularly crucial for time $t=N/2=30$, see the right panel in Fig. 4 where $m_{i}=10$, $t=30$ and the plot is with respect to the volume $R^{3}$.
Figure 4: The abrupt quench in the $(1+3)$-dimensional Minkowski space - spherical geometry; $N=60$, $m_{f}=0$. The left panel: $m_{i}=10$, $t=2$ slice; entropy (blue points) and capacity (red points) with respect to the area, $R^{2}$. The right panel: $m_{i}=10$, $t=30$ slice, the capacity with respect to the volume, $R^{3}$.
3.2 $(1+2)$ dimensions
For two spatial dimensions and constant mass, the numerical computations yield that the capacity like entropy linearly increases with the radius $R$ of the sphere (for the ground state of the discretized Hamiltonian), i.e. $S=a_{1}R$ and $C=a_{2}R$. However, as in three spatial dimensions the slopes ($a_{1}$ and $a_{2}$) for both of them are different and depend on mass. Using results from Sec. 2 and Appendix A, we plot the ratio $C/S$ (at the leading order, this is equal to $a_{2}/a_{1}$) with respect to $m$, see the left panel in Fig. 5. For the massless case this ratio is equal to $2.92$ and then increases with mass. Roughly, the values of this ratio are smaller than in the $(1+3)$-dimensional case, cf. Fig. 1, though for a larger $m$ the difference becomes negligible. We will return to this relation in Sec. 4 where the strip geometry in $1+2$ dimensions will be analysed.
Figure 5: The $(1+2)$-dimensional Minkowski space - the spherical geometry. The left panel: the ratio $C/S=a_{2}/a_{1}$ with respect to $m$; for $m=0$ it equals $2.92$. The right panel: the dynamics of the capacity for the abrupt quench, $m_{i}=10$, $m_{f}=0$, $N=60$; blue $n=10$, red $n=20$, green $n=30$ (vertical lines correspond to periods: $t=n$). The black vertical line corresponds to $t=N=60$ - a partial revival.
Now, let us consider the abrupt quenches. Similarly to the three-dimensional case entropy and capacity increase up to time $t=n$ and then the oscillations begin, see the right panel in Fig. 5. More precisely, after the initial growth, there is a “plateau" (with a very slow increase) which terminates about $t\simeq N-n$ and then, after $t\simeq N$, we have a revival of the entropy and capacity (see the black vertical line in Fig. 5). For further times the oscillations are around some asymptotic values. Such a behaviour can be interpreted in terms of the quasiparticles model presented in Ref. b5c for the $(1+1)$-dimensional case and finite-size integrable systems; we will analyse this issue in Sec. 4. Now, we will study the area law. Namely, using the decomposition analogous to the formula (16) (in the present case it contains $R$ and $R^{2}$ terms, respectively) we plot the ratio $b_{i}(t)/a_{i}(t)$ of the surface and volume terms in the $(1+2)$-dimensional case, see Fig. 6. Then, for sufficiently short time $t\ll R$ we observe the area law.
Figure 6: The ratio $b_{i}/a_{i}$ for the abrupt quench in the $(1+2)$-dimensional Minkowski space - spherical geometry; $N=60$, $m_{f}=0$. Blue curve - $b_{1}/a_{1}$ for entropy. Red curve - $b_{2}/a_{2}$ for capacity. The left panel: $m_{i}=10$. The right panel: $m_{i}=0.5$.
On the other hand, for larger times the area law does not hold. Namely, let us take $m_{i}=10$ as in the previous case. Then, from the left panel of Fig. 7 we infer that for a large initial mass the ratio $b_{i}(t)/a_{i}(t)$ oscillates, and the oscillations asymptotically settle down to some relatively large values (in our case about $0.5$ for the capacity and $0.2$ for entropy). Thus, for large times, the volume term is relevant. This can be also confirmed by plotting time slices for entropy and capacity as a function of $R^{2}$; in such a case we obtain quasi-linear behaviour.
Figure 7: The $(1+2)$-dimensional Minkowski space - the spherical geometry. Left panel: the abrupt quench, $m_{i}=10$, $m_{f}=0$, $N=60$, $n=30$; the blue curve - $b_{1}/a_{1}$ for entropy, the red curve - $b_{2}/a_{2}$ for capacity. The right panel: evolution of the capacity for the abrupt quench, $m_{i}=0.25$, $m_{f}=0$, $N=60$, $n=30$. The red line - theoretical values based on eqs. (17) and (21), the blue line the numerical results.
3.3 Theoretical predictions
Basing on the quasiparticles (EPR pairs) model, some theoretical results concerning the (Rényi) entropy for the abrupt quenches have been obtained in Refs. cotler ; casini . We will use these results to find a theoretical description of the dynamics of the capacity, and next compare them with the numerical computations. Namely, following the considerations of Ref. cotler , we obtain that the Rényi entropy in $1+2$ dimensions takes the form
$R_{\alpha}=s_{\alpha}\left\{\begin{array}[]{cc}2(t\sqrt{R^{2}-t^{2}}+R^{2}\arcsin(t/R))&t<R,\\ \pi R^{2}&t>R;\end{array}\right$. (17)
where
$s_{\alpha}=\frac{2\gamma_{E}+\psi(1-1/2\alpha)+\psi(1+1/2\alpha)+2\alpha(\ln 4-1)}{16\pi(\alpha-1)}m^{2}.$ (18)
while $\psi$ denotes digamma function. For $1+3$ dimensions we have
$R_{\alpha}=s_{\alpha}\left\{\begin{array}[]{cc}2\pi(R^{2}t-t^{3}/3)&t<R,\\ \frac{4\pi}{3}R^{3}&t>R;\end{array}\right$. (19)
where
$s_{\alpha}=\frac{4\alpha-3\cot(\pi/4\alpha)+\cot(3\pi/4\alpha)}{48\pi(\alpha-1)}m^{3}\,.$ (20)
Taking the limit $s=\lim_{\alpha\rightarrow 1}s_{\alpha}$ in eqs. (18) and (20) one gets $s=m^{2}\ln(2)/4\pi$ (in two spatial dimensions) and $s=m^{3}/(12\pi)$ (in three spatial dimensions), see Ref. cotler . Now, using formula (3) we get that the capacity $C$ has also the form of eqs. (17) and (19), respectively with the factor $s$ replaced by the constant $c$:
$c=\frac{\pi m^{3}}{16},\quad(1+3)\textrm{ dimensions};\quad\quad c=\frac{7\zeta(3)m^{2}}{16\pi}\simeq 0.167m^{2},\quad(1+2)\textrm{ dimensions.}$ (21)
Since the above formulae rely on a relatively simple model, they validity is limited and involves several assumptions, see Ref. cotler for more detailed discussion; in particular, the initial mass should be sufficiently small. Let us compare these models with our numerical results. First, according to the considerations from previous sections, we have the growth of capacity up to $t\simeq R$. More precisely, in $d=2$ dimensions the theoretical (see eq. (17) and below) and the numerical results are quite consistent when the initial mass is less than one. This can be seen especially for higher $n$; namely, taking $n=N/2$ and initial mass $m_{i}=0.5$ we see in the right panel of Fig. 7 that the theoretical and numerical results coincide quite well (the numerical plots are shifted to agree at $t=R$ ). Moreover, for initial times $t\ll R$ they yield the area law, what coincide very well with the numerical results presented in Sec. 3.2. For, the $(1+3)$-dimensional case the quasiparticles picture yield also the growth up to $t=R$ what coincides with the considerations obtained in Sec. 3.1. However, in this case the capacity (entropy) dynamics is not matched so well with the numerical results; this fact can be related to an additional contribution of the logarithmic divergence to the area law after quench in $1+3$ dimensions wilczek .
4 Minkowski spacetime - strip geometry
In this section we study the above issues for another geometry of entangling surface. This is interesting due to possible universal properties as well as validity of the quasiparticles model. To this end, following the reasoning of Ref. cotler , we can find the form of the capacity for the strip of width $l=2R$ in $1+d$ dimensions, when tracing over a $d$ dimensional slab of width $2R$ (the case $d=1$ corresponds to the interval, see b6l ). In such a case the system factorizes and the entropy as well as capacity reduce to the integral of their one dimensional counterparts, see Appendix A. Then using eq. (3) we readily get
$C_{strip}=\frac{A_{\perp}}{2^{d-2}\pi^{(d-1)/2}\Gamma(\frac{d-1}{2})}\int_{0}^{\infty}dkk^{d-2}C(R,m^{2}(t)+k^{2}),$ (22)
where $A_{\perp}$ is the width of the strip in the perpendicular direction and $C$ is the capacity for the one dimensional system (i.e. for the interval of the length $2R$).
On the other hand, for $d=2$ and the periodic boundary conditions the quasiparticles model cotler ; casini implies the following form of the capacity
$C_{strip}^{t}=\frac{4A_{\perp}c}{\pi}\left\{\begin{array}[]{cc}t&t<R,\\ t-\sqrt{t^{2}-R^{2}}+R\arccos(R/t)&t>R,\end{array}\right$. (23)
where $c$ is given by eq. (21). For the Dirichlet boundary conditions we should make the replacement $c\rightarrow c/2$ and the change is at $t=2R$. In what follows we will compare the theoretical model (23) with the numerical results based on eq. (22) as well as with the previous ones for spheres.
First, let us analyse the case of constant mass. To this end we compute the entropy and capacity with respect to $n$ (equivalently $R$). To fix attention we present results in $d=3$ spatial dimensions, see the left panel of Fig. 8; for the two spatial dimensions the results are similar. Namely, we observe that for sufficient large radius $R$ both quantities are almost constant (here we present the case $m=0$; however, for other values of $m$ this is even more evident) and thus do not depend on $R$. In consequence, only the transversal area remains and the area law holds. In view of this the ratio $C/S$ is constant, i.e. it depends on $m$ only (at the leading order). Let us compare this ratio with the one obtained for the spherical geometry. Namely, in the right panel of Fig. 8 we present this ratio as a function of mass (here we consider $(1+3)$-dimensional case, but a similar situation holds for two spatial dimensions). Comparing the right panels of Fig. 1 and Fig. 8 we see that $C/S$ is the same for both geometries with very good accuracy (the same holds also in two spatial dimensions). Thus despite a quite different geometry of the entangling surface, $C/S$ does not change. This observation suggests certain universality of this ratio.
Figure 8: The $(1+3)$-dimensional Minkowski space - the strip geometry, $N=60$. The left panel $m=0$, entropy - blue points, capacity - red points, with respect to $n$. The right panel: the ratio $C/S$ (for higher $R$) with respect to $m$, cf. Fig. 1 for the spherical geometry.
Now let us consider the quenches for the strip geometry. Using eq. (22) we can compute the dynamics of the capacity. In two spatial dimensions the numerical results for the initial times are presented, for $n=30$, by the red curve in Fig. 9. We observe, in particular, a linear growth up to $t=n=2R=30$ for the Dirichlet (and $t=n/2=15$ periodic, respectively) boundary conditions. More importantly, with an appropriate choice of initial mass (i.e. about $m_{i}=0.5$) the slope agrees very well with the theoretical predictions following from the formula (23), see the red curve in Fig. 9 (as usual we shift the plots to match them). For further time ($R\ll t$ ), similar to the spherical geometry, the oscillatory behaviour appears, see the left plot in Fig. 10. In view of this let us study the role of the volume term. It turns out that, similarly to the spherical geometry, for a larger $m_{i}$ the contribution related to the volume factor becomes more significant; this can be especially seen by considering the time slices for large times. Namely, taking $m_{i}=10$ the numerical computations give that the capacity increases linearly with $n$ (at least for suitable radius). For three dimensional case we observe again the linear growth at the initial times; however, the slope (except small masses) is different from the one obtained by means of the quasiparticles model (similarly to the spherical case).
Figure 9: The $(1+2)$-dimensional Minkowski space - the strip geometry. The dynamics of the capacity for the abrupt quench $m_{i}=0.25$, $m_{f}=0$, $N=100$, $n=30$. The theoretical capacity - the red curve, the numerical results - green points. The left panel: the Dirichlet boundary conditions. The right panel: the periodic boundary conditions.
Now, let us recall that for the spherical case, see Fig. 5, we have some plateaus and revival times in the entanglement dynamics; here, see the left panel in Fig. 10, we observe an analogous situation (for the Dirichlet boundary conditions the first revival time corresponds to $t=N=100$). To gain some insight into this issue let us plot the evolution of the entanglement entropy (a similar situation holds for the entanglement capacity) for different values of $N$ and the periodic boundary conditions. Namely, in the right panel of Fig. 10 we present the cases $N=100$ and $N=200$ with $n=20$. Then, after the linear growth, for $n/2\lesssim t\lesssim(N-n)/2$ we observe a plateau with slow saturation. Next, the plateau terminates at $t\simeq(N-n)/2$ (the first quasiparticles produced at the boundary of the subsystem re-enter due to periodic boundary conditions). Such a process lasts until $t\simeq N/2$ when we observe the entanglement revival and the dynamics restarts. In view of this, the quasiparticles mechanism proposed in Ref. b5c for finite-size systems can be applied also in higher dimensions (in our case, due to the periodic boundary conditions, the final mass is $m_{f}=0.01$ and thus the maximum quasiparticle velocity is almost one, the speed of light).
Figure 10: The quench in the $(1+2)$-dimensional Minkowski space - the strip geometry. The left panel: the capacity evolution for further times, $m_{i}=0.25$, $m_{f}=0$, $N=100$, the Dirichlet boundary conditions. Blue - $n=10$, green - $n=20$, red - $n=50$. The right panel: the entropy dynamics for the periodic boundary conditions ($m_{i}=1$ and $m_{f}=0.01$) for $n=20$; $N=100$ the red dashed line, $N=200$ the green line. The black vertical lines denote the revival times following from the quasiparticles model (i.e. $n=N/2$).
5 Universe expansion
In this section we study the scalar field $\Phi$ in the curved spacetime. More precisely, we consider the FLRW metric
$ds^{2}=dt^{2}-a^{2}(t)d{\bf{x}}^{2}=a^{2}(\eta)(d\eta^{2}-d{\bf{x}}^{2}),$ (24)
in the cosmic time $t$; or alternatively, in the conformal time $\eta$, $dt=a(\eta)d\eta$.
Let us start with the $(1+1)$-dimensional case and conformal time. Then, following the standard discretization procedure (with lattice spacing equals one) we arrive at the Hamiltonian
$H(\eta)=\frac{1}{2}\sum_{j}\left(\pi_{j}^{2}+(\phi_{j}-\phi_{j+1})^{2}+m^{2}(\eta)a^{2}(\eta)\phi_{j}^{2}\right).$ (25)
Let us note that the Hamiltonian (25) under the Dirichlet boundary condition can be written in the form (2) with $\Lambda(\eta)$ given by eq. (38) with $M^{2}=m^{2}(\eta)a^{2}(\eta)$. Thus the eigenvalues of $\Lambda(\eta)$ read
$\lambda_{j}(\eta)=\lambda_{j}^{0}+m^{2}(\eta)a^{2}(\eta),$ (26)
where $\lambda^{0}_{j}$ are the (constant) eigenvalues of the matrix (38) with $M=0$, i.e. $\lambda_{j}^{0}=4\sin^{2}(j\pi/N)$, $j=1,\ldots,N$.
Alternatively, for the cosmic time we get
$\hat{H}(t)=\frac{1}{2}\sum_{j}\left(\hat{\pi}_{j}^{2}+\frac{(\hat{\phi}_{j}-\hat{\phi}_{j+1})^{2}}{a^{2}(t)}+\hat{\Omega}(t)\hat{\phi}_{j}^{2}\right),$ (27)
where $\hat{\Omega}(t)=m^{2}(t)+\dot{a}^{2}/4a^{2}-\overset{..}{a}/2a$. The equivalence of the descriptions in both times can be confirmed by the following canonical time-dependent transformation
$\phi_{j}=\hat{\phi}_{j}/\sqrt{a},\quad p_{j}=\sqrt{a}\hat{p}_{j}-\dot{a}\hat{\phi}_{j}/(2\sqrt{a})\ .$ (28)
In fact, we have the identity
$H(\eta(t))\frac{d\eta}{dt}+\frac{\partial F}{\partial t}=\hat{H}(t),$ (29)
where $F$ is the generating function of the transformation (28), i.e.
$F(\phi_{1},\ldots,\phi_{N},\hat{p}_{1},\ldots,\hat{p}_{N},t)=\sum_{j}(\sqrt{a}\phi_{j}\hat{p}_{j}-\dot{a}\phi_{j}\phi_{j}/4).$ (30)
The above transformation implies that both descriptions are equivalent; in consequence the symplectic covariance argument, see Ref. sitter , implies that the entropy and capacity of entanglement are equivalent for both realizations. Below, we will show this explicitly by considering the evolution of the initial ground state, see Sec. 2. Namely, let us take the functions $b_{j}(\eta)$ defining the evolution of the state in the conformal time, i.e. satisfying the Ermakov equation (5) with $\lambda_{j}(\eta)=m^{2}(\eta)a^{2}(\eta)+\lambda_{j}^{0}$ . Then, we can readily check that the functions
$\hat{b}_{j}(t)=b_{j}(\eta(t))\sqrt{a(t)},$ (31)
satisfy the Ermakov equation in the cosmic time (with $\hat{\lambda}_{j}(t)=\hat{\Omega}(t)+\lambda^{0}_{j}/a^{2}(t)$). This together with the results presented in Sec. 2 yield $\hat{\tilde{B}}=(\tilde{B}+\dot{a}I/4)/a$, thus $\hat{\tilde{B}}_{2}=\tilde{B}_{2}/a$; on the other hand, we have $\hat{\Omega}=\Omega/a$. In consequence, by virtue of eq. (10) the eigenvalues of the matrix (13) (and thus (12)) do not change and consequently the Rényi entropies as well (we have to make only the substation $S(\eta(t))=\hat{S}(t)$). In particular, when $m=0$ then $b_{j}=1$ and thus $S(\eta)=const$. On the other hand, we have $\hat{b}_{j}=\sqrt{a}$ and the matrix $\hat{B}$ is diagonal; in consequence, $\hat{B}_{2}=0$ and then the matrix $\hat{\Upsilon}$ and $\hat{\Delta}$ contain the time dependency through the same common factor $1/a$; consequently, $\hat{\Delta}$ is time-independent and $\hat{S}$ and $\hat{C}$ are also constant.
A similar situation holds in $1+3$ dimensions. Namely, for the conformal time the discretization procedure (enhanced by the simple canonical transformation i.e. the scaling of momenta by $a$ and coordinates by $1/a$) yields that the Hamiltonian $H$ is the sum $H(\eta)=\sum_{lm}H_{lm}(\eta)$ where
$H_{lm}(\eta)=\frac{1}{2}\sum_{j}\left(\pi_{lm,j}^{2}+(j+\frac{1}{2})^{2}\left(\frac{\phi_{lm,j}}{j}-\frac{\phi_{lm,j+1}}{j+1}\right)^{2}+\frac{l(l+1)}{j^{2}}\phi^{2}_{lm,j}+\Omega(\eta)\phi_{lm,j}^{2}\right),$ (32)
with $\Omega(\eta)=m^{2}(\eta)a^{2}(\eta)-\frac{a^{\prime\prime}(\eta)}{a(\eta)}$. Thus, the Hamiltonian $H_{lm}(\eta)$ takes the form (2) with $\Lambda$ given by eq. (40) with $M^{2}=\Omega(\eta)$. In consequence, the eigenvalues are of the form $\lambda_{j}=\lambda_{j}^{0}+\Omega(\eta)$ where $\lambda_{j}^{0}$ are the (constant) eigenvalues of the matrix (40) with $M=0$. On the other hand, performing the canonical transformation (28) we obtain the Hamiltonian in the cosmic time $\hat{H}(t)=\sum_{lm}\hat{H}_{lm}(t)$.
$\hat{H}_{lm}(t)=\frac{1}{2}\sum_{j}\left(\hat{\pi}_{lm,j}^{2}+\frac{(j+\frac{1}{2})^{2}}{a^{2}(t)}\left(\frac{\hat{\phi}_{lm,j}}{j}-\frac{\hat{\phi}_{lm,j+1}}{j+1}\right)^{2}+\frac{l(l+1)}{j^{2}a^{2}(t)}\hat{\phi}^{2}_{lm,j}+\hat{\Omega}(t))\hat{\phi}_{lm,j}^{2}\right),$ (33)
where $\hat{\Omega}(t)=m^{2}(t)-\frac{3\dot{a}^{2}(t)}{4a^{2}(t)}-\frac{3\overset{..}{a}(t)}{2a(t)}$. As above the dynamics of the initial state governed by both the Hamiltonians is determined by the functions $b_{j}(\eta)$ and $\hat{b}_{j}(t)$ respectively, which satisfy the suitable Ermakov equations (with the frequencies $\lambda_{j}$ and $\hat{\lambda}_{j}$, containing $\Omega$ and $\hat{\Omega}$, respectively). Now, by straightforward calculations we check that the relation (29) holds. In consequence, the suitable entropies coincide in both pictures.
5.1 De Sitter space
For dS space we have $a(t)=e^{Ht}$, equivalently $a(\eta)=(-1)/H\eta$ for $\eta<0$. The dynamics of the entanglement entropy for dS space (and the standard vacuum states) was studied in Ref. sitter2 and more recently in Ref. sitter0 for a field at two distinct spatial locations sitter0b , as well as in the lattice approach in Refs. sitter ; sitter3 ; sitter4 . In particular, it has been argued therein that at the leading order the area law holds with respect to the proper area of the surface; for the comoving coordinates the suitable $\eta$-dependence arises. In what follows we assume the Bunch-Davies (BD) vacuum, then the solutions of the Ermakov equations, i.e. $b$’s functions, tend to one while their derivatives tend to zero for $\eta\rightarrow$-$\infty$. In consequence, they read
$b_{j}^{2}(\eta)=-\frac{\pi}{2}\sqrt{\lambda_{j}^{0}}\eta\left(J_{\nu}^{2}(-\eta\sqrt{\lambda_{j}^{0}})+Y^{2}_{\nu}(-\eta\sqrt{\lambda_{j}^{0}})\right),$ (34)
where, $J_{\nu}$ and $Y_{\nu}$ are the Bessel functions, while, in $1+3$ dimensions, $\lambda^{0}_{j}$ are the eigenvalues of the matrix (40) with $M=0$, while $\nu=\sqrt{9-4m^{2}/H^{2}}/2$. The typical dynamics of the capacity is presented in the left panel of Fig. 11 (for $n=20$ and n=30n=30) [^4] To make contact with our previous considerations, we follow the regularization procedure of counting $l$ from previous sections. However, a different regularization can also be used, see Ref. cosmo , then subhorizon modes are excluded and $l$ is truncated earlier.. Similarly to the entropy, the capacity increases when $\eta$ approaches to zero; moreover, it increases with $n$ (radius). To analyse the latter issue and the area law, we plot time slices with respect to $n$. Then, we observe in the right panel of Fig. 11, that the area law holds also for the capacity. Namely, for a fixed time (even small $\eta=-0.2$) the values fit very well into the parabola.
Figure 11: $(1+3)$-dimensional dS space, $N=60$, $m=0$, $H=1$. The left panel: capacity $n=10$ (yellow), $n=20$ (red). The right panel: the time-fixed slice $\eta=-0.2$ (together with the parabolic fitting) - entropy (blue) and capacity (red).
In view of this the ratio of the capacity and entropy (at the leading order) does depend on the radius. In previous sections for massless field in the $(1+3)$-dimensional Minkowski spacetime we obtained that this ratio is equal to $5.2$; now, let us analyse this problem for dS space. To this end we plot the ratio $C/S$ for several values of $n$, see Fig. 12.
Figure 12: $(1+3)$-dimensional dS space. $N=60$, $m=0$, $H=1$. The ratio of the capacity and entropy for $n=10$, (green) $n=20$ (blue), and $n=30$ (red).
First, we observer that the ratio indeed does not depend on $n$ ($n=10,20,30$ coincide); moreover, it is constant and approximately equals five for initial times (this is in agreement with the considerations presented in Sec. 3 and the definition of the BD state). However, as $\eta$ approaches to zero $C/S$ is decreasing. The numerical results, see the right plot in Fig. 12, yields that this ratio tends to one in the limit $\eta\rightarrow 0^{-}$ :
$\frac{C}{S}\simeq 5,\ \textrm{ for }\eta\rightarrow-\infty,\quad\frac{C}{S}\simeq 1,\ \textrm{ for }\eta\rightarrow 0^{-}\ .$ (35)
According to the results obtained in Ref. boer , the above observation may suggest that the scalar field theory in dS space at late times is related to conformal field theories (despite the fact that in full dS space we do not have the conformal covariance, i.e. in general $m^{2}\neq 2H^{2}$). This idea fits into the recent investigations of dS/CFT correspondence latetime1 ; latetime2 ; latetime3 , where the behaviour of the scalar fields in dS space at late times (i.e. $\eta\rightarrow 0^{-}$) is associated with certain (Euclidean) conformal field theories (more precisely, generating functional for correlation functions of the dual conformal theories). Although the description of this correspondence is not complete currently, following the AdS/CFT correspondence, a further analysis of the entanglement structure at late times in dS space would be valuable.
5.2 Radiation-dominated era
Now, we let us consider other FLRW metrics. Namely, we will study the transition from dS space to the radiation-dominated era; the latter will be modeled by a metric with linear function $a(\eta)\sim\eta$ (by similar considerations we can add the era of matter domination, $a(\eta)$ is a quadratic function). In such a case, $a^{\prime\prime}(\eta)=0$ and thus for the massless field $\Omega=0$; then, in turn, the function $b_{j}$ can be readily found. It turns out that, after the transition the area law breaks and a volume term develops giving contribution to the entropy at late times cosmo ; cosmo2 . A similar situation appears for the capacity. Namely, assuming the transition from dS to the RD era takes place at $\eta=-1$, we observe that the monotonic growth is broken and the quasi-periodicity appears during the evolution, see the left panel in Fig. 13 for the initial times. Moreover, after transition the area law does not hold (the contribution from the volume term appears), see dots in the right panel in Fig. 13 where the entropy and capacity for $\eta=50$ are depicted together with suitable parabolas.
Figure 13: The transition from $(1+3)$-dimensional dS space to the radiation-dominated era (at $\eta=-1$); $N=60$, $m=0$, $H=1$. The left panel: evolution of the capacity $n=20$ (black), $n=30$ (blue). The right panel: time-fixed slice for $\eta=50>0$ and the parabolic fitting - entropy (blue) and capacity (red).
To analyse this situation in more detail we split the dynamics of the capacity into two parts: $a_{2}$ and $b_{2}$, related to the quadratic and cubic parts, see eq. (16). The temporal evolution of these coefficients are presented in Fig. 14. We observe that at the beginning the area law holds with good approximation; however, for further times the cubic term develops, and the situation repeats quasi-periodically; finally, the oscillations of the cubic part decay with time and asymptotically settle to a constant value (in our case $b_{2}\simeq 0.3$).
Figure 14: The transition from $(1+3)$-dimensional dS space to the radiation-dominated era (at $\eta=-1$); $N=60$, $m=0$, $H=1$. The temporal evolution of the capacity. The left plot: the square factor $a_{2}$. The right plot: the cubic factor $b_{2}$, cf. eq. (16).
5.3 Quenches in de Sitter space
Now, we will follow Sec. 3 and consider quench phenomena in dS space. This is interesting due to the fact that rapid phenomena during the early expansion of the universe (e.g. inflation, reheating, particles production) should be related to the non-equilibrium processes, see e.g. cosmo3 ; cosmo4 ; cosmo5 and references therein. To get some insight into these issues, in particular thermalisation (equilibration), quenches in closed systems are frequently considered. In view of this, we will study the evolution of entanglement entropy and capacity by means of an abrupt quench in dS space; with the special attention put on the transition from the area law to volume law (the latter is characteristic for thermal entropy).
To this end, let us recall that the mass quench in the FLRW spacetime is described by the Hamiltonian (32) with effective mass $\Omega(\eta)=m^{2}(\eta)a^{2}(\eta)-a^{\prime\prime}(\eta)/a(\eta)$. Thus, the mass quench in the FLRW space can be recast as the mass quench (relatively to effective mass $\Omega$) in the ordinary Minkowski spacetime (if we add conformal coupling term $\xi R$ to the action, then we get a more direct expression $\Omega(\eta)=m^{2}(\eta)a^{2}(\eta)$). Consequently, in this picture, the abrupt quench in the FLRW space (in particular for dS space) does not give an instantaneous cooling in the Minkowski spacetime needed for the quasiparticles model. Moreover, as mentioned in Sec. 3, for the Minkowski spacetime in three spatial dimensions we encounter some additional problems; for dS space the situation becomes more complicated, i.e. the entropy contains additional terms related to the Hubble constant (in particular the long range contribution to the entanglement entropy). In view of this, to analyse the entanglement dynamics, we follow the lattice approach presented in previous sections. To this end we need the functions $b_{j}$ describing the dynamics of the reduced density.
More specifically, let us assume that we start with a massive field $m_{i}$ and next there is an abrupt change of the mass parameter to zero value (the massless field). Then, the frequencies appearing in the Ermakov equation are give by the formula
$\lambda_{j}(\eta)=\lambda_{j}^{0}+\left\{\begin{array}[]{c}(\frac{m_{i}^{2}}{H^{2}}-2)\frac{1}{\eta^{2}},\quad\eta<\eta_{0},\\ -\frac{2}{\eta^{2}}\quad\eta\in[\eta_{0},0),\end{array}\right$. (36)
where $\eta_{0}<0$ is a fixed point. Assuming the BD vacuum state the functions $b_{j}$ are described by eq. (34) for $\eta<\eta_{0}$ . It remains to find $b_{j}$ for $\eta\in[\eta_{0},0)$ in such a way that they as well their derivatives are continuous at the point $\eta_{0}$. This can be done by the straightforward but rather tedious computations. The final result reads
$b_{j}(\eta)=\sqrt{x_{j}^{2}(\eta)+y_{j}^{2}(\eta)/A^{2}},\quad\eta\in[\eta_{0},0),$ (37)
where the functions $x_{j}(\eta),y_{j}(\eta)$ and the constant $A$ are given in Appendix B. Now, we are in the position to analyse the dynamics of the entropy and capacity. Of course, for $\eta<\eta_{0}$, we have the monotonically increasing growth of the entropy and capacity. Since the frequencies $\lambda_{j}$ in both cases contain the factor $1/\eta^{2}$ we expect also a similar behaviour for $\eta\rightarrow 0$. However, for the intermediate times (related to the initial mass) this situation may change. Indeed, we observe in Fig. 15 that after the change of mass (here, from $m_{i}=\sqrt{5}/2$ to zero) at $\eta_{0}=-10$, there is a period of time resembling quasi-oscillatory behaviour in the Minkowski spacetime, and next (for $\eta>-7$) again both quantities uniformly increase.
Figure 15: The quench in $(1+3)$-dimensional dS space; $N=60$, $H=1$, the mas quench at $\eta_{0}=-10$, the initial mass $m_{i}=\sqrt{5}/2$ and $n=30$. The left plot: entropy, the right plot: capacity.
However, it seems that there is a one difference to the Minkowski case; namely, for these intermediate times the area law is preserved with good approximation. Indeed, for both the entropy and capacity the coefficients corresponding to the cubic terms are very small compared to the quadratic terms (i.e. $b_{i}/a_{i}\simeq 0.001$ for $i=1,2$); in consequence, the leading contribution comes from the quadratic term, see Fig. 16, where the evolution of the quadratic factor $a_{2}$ for the capacity is presented (cf. the right plot in Fig. 15) as well as a time fixed slice together with the parabola approximation for $\eta=-9$ (an intermediate time). Consequently, the area law holds for all times with good accuracy. This situation is different from that obtained in Sec. 3, where the volume term after quench was significant. Similarly to the free theory in the Minkowski spacetime (in general, for the integrable models), we expect that the thermalisation takes place in a more general sense, i.e. the steady state is described by a generalized Gibbs ensemble (due to the constrains related to the conserved quantities), see e.g. gge ; gge2 and references therein. However, in dS space the absence of volume term implies further complications; the injection of energy does not essentially change the entropy, in other words, entropy attains the maximum value; consequently, it is related to a uniform probability distribution. This, in turn, suggest that thermalisation takes place for (large) infinite temperature. Although, such a reasoning is quite superficial, it has some points in common with recent ideas concerning thermodynamic of dS space presented in Ref. dst . To get more deeper insight into the above issue and to better understand non-equilibrium processes in dS space further studies (including continuous quenches, interacting fields and the symmetry breaking mechanisms) are needed.
Figure 16: The quench in $(1+3)$-dimensional dS space; $N=60$, $H=1$, the mass quench at $\eta_{0}=-10$, the initial mass $m_{0}=\sqrt{5}/2$. The left plot: the quadratic factor $a_{2}$ for the capacity. The right panel: entropy (blue) and capacity (red) – the time-fixed slice (at $\eta=-9>\eta_{0}$) after quench (together with the parabolic fitting).
6 Conclusions
In this work, we have studied some aspects of entanglement in quantum field theory; in particular, the ones related to the area law. To get more insight into these issues, besides entropy, we use the notion of the capacity of entanglement which gained recently some attention and can be treated as a measure of entropy fluctuations. Both the quantities together can provide more information on the entanglement structure and can be helpful in finding some universal properties of the theory. In our investigations, we have considered physically more interesting, but less commonly studied, higher dimensional spacetimes including some curved backgrounds (relevant for cosmology). Moreover, the special attention has been put on quench phenomena, which are useful in various physics contexts (such as thermalisation processes, phase transitions and the physics of early universe). In the latter case, the time-dependent mass parameter has been considered and the numerical results were compared with the dynamics resulting from theoretical models. Let us now summarize our results.
First, we showed that for the fields with constant mass, the capacity, like entropy, exhibits the area law (at the leading order). We observe this for two kinds of geometries of the entangling surface in the Minkowski spacetime: spheres and strips. This observation implies that the ratio of both quantities does not depend on the area, and more importantly, this ratio takes the same values for both kinds of geometries. Next, we turned our attention to quenches. First, we analysed the dynamics of capacity and showed that after some initial time the area law is broken and the volume term comes into play; for sufficiently strong abrupt quenches this term can be crucial. Moreover, for the strip geometry the initial growth is linear and does not depend on the width of the strip (thus the area law holds). To get some insight into this issue we compared these results with theoretical predictions resulting from the quasiparticles model; in $1+2$ dimensions we got good agreement for both geometries (including the revivals times for the entanglement dynamics).
In the second part we considered the above issues in curved spacetimes. We started by showing explicitly that for fields in the FLRW space the description of the entropy (capacity) in the cosmic and conformal times are equivalent. Next, we specialized the metric taking dS space, and then the metric modeling transition to the radiation-dominated era. For the BD state of dS space, similarly to the entropy, the area scaling holds also for the capacity. Thus the ratio $C/S$ does not depend on the radius and it tends to one as conformal time approaches zero; the latter behaviour is interesting due to the recent investigations of dS/CFT correspondence. Moreover, we showed that after the transition to the radiation-dominated era the volume term develops in the dynamics of the capacity. Finally, we analysed the evolution of the entropy and capacity during the abrupt quench in dS space. First, we found the functions describing the evolution of the state in such a scenario. Next, using these results, we showed that after quench the area law survives with good accuracy.
Of course, the above results do not exhaust the subject. In this context it would be interesting to consider continuous or multiple quenches b12d , higher derivative theories b20a or different vacuum states in dS space b20b . The notion of the modular entropy b20c and boundary quenches can be also examined cotler . Finally, following Refs. b20d ; b20e the consequences of non-minimal coupling terms and regular black holes are also worth of study.
Acknowledgements.
The author would like to thank Mohammad Reza Mohammadi Mozaffar for valuable discussion and references as well as Piotr Kosiński for useful remarks. Special thanks are due to anonymous referee for valuable comments and suggestions.
Appendix A Discretization procedure
In this appendix, we briefly recall the form of the discretized Hamiltonians and the corresponding entropies for two geometries of the entanglement surface in the Minkowski spacetime, see e.g. Refs. cotler ; b13a ; b13b ; b13c . These facts turn out to be useful for the FLRW metrics presented in Sec. 5.
Let us consider the scalar field with mass $M$. In 1+1 dimensions (and the Dirichlet boundary conditions) the matrix $\Lambda$ in eq. (2) is of the form
$\Lambda_{jj}=2+M^{2},\quad\Lambda_{j,j+1}=\Lambda_{j+1,j}=-1.$ (38)
In 1+2 dimensions the discretized Hamiltonian is the sum $H=\sum_{l=-\infty}^{\infty}H^{l}$ where $H^{l}$ are of the form (2) with the following $\Lambda^{l}$
$\Lambda^{l}_{11}=\frac{3}{2}+l^{2}+M^{2},\quad\Lambda^{l}_{jj}=2+\frac{l^{2}}{j^{2}}+M^{2},\quad\Lambda^{l}_{j,j+1}=\Lambda^{l}_{j+1,j}=-\frac{(j+1/2)}{\sqrt{j(j+1)}}.$ (39)
In view of this the Rényi entropy for the Gaussian state is the sum of the $l$ components $R_{\alpha}=R_{\alpha}^{0}+2\sum_{l=1}^{\infty}R_{\alpha}^{l}$; in consequence, the same holds for the entropy and capacity.
For $1+3$ dimensions and the spherical geometry we have $H=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}H^{lm}$ where $H^{lm}$ is described by (2) with
$\Lambda^{lm}_{11}=\frac{9}{4}+l(l+1)+M^{2},\quad\Lambda^{lm}_{jj}=2+\frac{1}{(2j^{2})}+\frac{l(l+1)}{j^{2}}+M^{2},\quad\Lambda^{lm}_{j,j+1}=\Lambda^{lm}_{j+1,j}=-\frac{(j+1/2)^{2}}{j(j+1)}.$ (40)
Thus $R_{\alpha}=\sum_{l=0}^{\infty}(2l+1)R_{\alpha}^{l}$ and analogously for the entanglement entropy and capacity.
For the strip geometry in the $(1+d)$-dimensional Minkowski spacetime we trace over a $d$-dimensional slab of width $2R$ and the cross-sectional area $A_{\perp}$. Then, see cotler , the Hamiltonian factorizes
$H=\frac{A_{\perp}}{(2\pi)^{d-1}}\int d^{d-1}k_{\perp}\tilde{H}(q_{k_{\perp}},p_{k_{\perp}},M^{2}+k_{\perp}^{2}),$ (41)
where $\tilde{H}$ is the Hamiltonian of a massive (with the mass parameter $M^{2}+k_{\perp}^{2}$) field in $1+1$ dimensions; $k_{\perp}$ denotes momentum in the perpendicular direction. The modes $k_{\perp}$ decouple and, consequently, the Rényi entropy can be reduced to the integral over one-dimensional counterparts. This together with formula (3) give the capacity described by eq. (22).
Appendix B Quench of the Bunch-Davies vacuum
In this appendix we compute the solutions of the Ermakov equations describing the quench of the BD state in $(1+3)$-dimensional dS space. Namely, let us assume that mass changes at $\eta=\eta_{0}<0$ from $m_{i}$ to zero, and for $\eta<\eta_{0}$ we have the BD vacuum state. Then $\lambda_{j}(\eta)$ are given by eq. (36) where $\lambda^{0}_{j}$ are the eigenvalues of the matrix (40) with $M=0$. Moreover, for $\eta\leq\eta_{0}$ the solutions of the Ermakov equations (5) are given by formula (34). Now, we will find $b$’s after quench (demanding that their derivatives are continuous at $\eta_{0}$). After straightforward but tedious computations, we get that the functions $b_{j}(\eta)$ for $\eta\in[\eta_{0},0)$ are given by eq. (37) with
$x_{j}(\eta)=C_{j}\left(\frac{\sin(\sqrt{\lambda_{j}^{0}}\eta)}{\sqrt{\lambda_{j}^{0}}\eta}-\cos(\sqrt{\lambda_{j}^{0}}\eta)\right)+D_{j}\left(\sin(\sqrt{\lambda_{j}^{0}}\eta)+\frac{\cos(\sqrt{\lambda_{j}^{0}}\eta)}{\sqrt{\lambda_{j}^{0}}\eta}\right),$ (42)
$\begin{split}y_{j}(\eta)=&\left(\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})+\frac{\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})}{\sqrt{\lambda_{j}^{0}}\eta_{0}}\right)\cdot\left(\cos(\sqrt{\lambda_{j}^{0}}\eta)-\frac{\sin(\sqrt{\lambda_{j}^{0}}\eta)}{\sqrt{\lambda_{j}^{0}}\eta}\right)+\\ &\left(\frac{\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})}{\sqrt{\lambda_{j}^{0}}\eta_{0}}-\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})\right)\cdot\left(\sin(\sqrt{\lambda_{j}^{0}}\eta)+\frac{\cos(\sqrt{\lambda_{j}^{0}}\eta)}{\sqrt{\lambda_{j}^{0}}\eta}\right),\end{split}$ (43)
where
$\begin{split}C_{j}=&\frac{A_{j}}{\lambda_{j}^{0}}\left(\frac{\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})}{\eta_{0}^{2}}+\frac{\sqrt{\lambda_{j}^{0}}}{\eta_{0}}\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})-\lambda_{j}^{0}\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})\right)+\\ &\frac{B_{j}}{\lambda_{j}^{0}}\left(\sqrt{\lambda_{j}^{0}}\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})+\frac{\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})}{\eta_{0}}\right),\end{split}$ (44)
$\begin{split}D_{j}=&\frac{A_{j}}{\lambda_{j}^{0}}\left(\lambda_{j}^{0}\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})-\frac{\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})}{\eta_{0}^{2}}+\frac{\sqrt{\lambda_{j}^{0}}}{\eta_{0}}\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})\right)+\\ &\frac{B_{j}}{\lambda_{j}^{0}}\left(\sqrt{\lambda_{j}^{0}}\cos(\sqrt{\lambda_{j}^{0}}\eta_{0})-\frac{\sin(\sqrt{\lambda_{j}^{0}}\eta_{0})}{\eta_{0}}\right),\end{split}$ (45)
and $b_{j}(\eta_{0})=A_{j}$, $\frac{db_{j}}{d\eta}(\eta_{0})=B_{j}$ are the values for the BD state at $\eta=\eta_{0}$ (see eq. (34)).